# PriorRG: Prior-Guided Contrastive Pre-training and Coarse-to-Fine Decoding for Chest X-ray Report Generation

### v1

@@ Abstract
Chest X-ray report generation aims to reduce radiologists’ workload by automatically producing high-quality preliminary reports. A critical yet underexplored aspect of this task is the effective use of patient-specific prior knowledge—including clinical context (e.g., symptoms, medical history) and the most recent prior image—which radiologists routinely rely on for diagnostic reasoning. Most existing methods generate reports from single images, neglecting this essential prior information and thus failing to capture diagnostic intent or disease progression. To bridge this gap, we propose PriorRG, a novel chest X-ray report generation framework that emulates real-world clinical workflows via a two-stage training pipeline. In Stage 1, we introduce a prior-guided contrastive pre-training scheme that leverages clinical context to guide spatiotemporal feature extraction, allowing the model to align more closely with the intrinsic spatiotemporal semantics in radiology reports. In Stage 2, we present a prior-aware coarse-to-fine decoding for report generation that progressively integrates patient-specific prior knowledge with the vision encoder’s hidden states. This decoding allows the model to align with diagnostic focus and track disease progression, thereby enhancing the clinical accuracy and fluency of the generated reports. Extensive experiments on MIMIC-CXR and MIMIC-ABN datasets demonstrate that PriorRG outperforms state-of-the-art methods, achieving a 3.6% BLEU-4 and 3.8% F1 score improvement on MIMIC-CXR, and a 5.9% BLEU-1 gain on MIMIC-ABN. Code and checkpoints will be released upon acceptance.
@@ 1 Introduction
Figure 1: (a-c) show the workflow of existing methods and our approach. (d) summarizes their key properties. Longitudinal data includes both current and prior images. Patient-specific prior knowledge comprises the prior image, indication, and medical history, which may be partially missing. “CC” and “DP” indicate whether a method models clinical context and disease progression.
@@
Radiology report generation (RRG) (Wang et al. 2025b; Mei et al. 2024) leverages AI techniques to automatically interpret medical images—such as chest X-rays (Hyland et al. 2024; Zhang et al. 2025; Nowak et al. 2025; Wang et al. 2025a), CT scans (Hamamci, Er, and Menze 2024; Li et al. 2025a), and pathological slides (Guo et al. 2024; Chen et al. 2024)—and generate structured textual descriptions of clinically relevant findings. This automation supports radiologists by providing standardized preliminary reports, thereby improving diagnostic efficiency and consistency.
@@
While recent RRG approaches have achieved remarkable progress, most models operate on isolated images (see Figure 1(a)) and overlook essential patient-specific prior knowledge, including both clinical context (i.e., the indication and history sections) and the most recent prior image. Such prior knowledge plays a critical role in real-world clinical reasoning by enabling personalized interpretation and tracking of disease progression. However, existing methods (Li et al. 2023; Liu et al. 2024a; Xiao et al. 2025) largely ignore these factors, limiting their ability to produce context-aware, disease progression-oriented reports.
@@
To capture the clinical context, (Nguyen et al. 2023) treats indications as auxiliary input while ignoring the noise in the data. SEI (Liu et al. 2024b) eliminates invalid or corrupted characters via preprocessing and employs a cross-modal fusion network to incorporate indications, thereby generating accurate findings. Similarly, for (Mei et al. 2025). However, they fail to consider longitudinal information, often resulting in hallucinations when describing disease progression. To combat this issue, existing methods incorporate prior images (Figure 1(b)) and employ techniques such as report pre-filling (Zhu et al. 2023), intra-modality similarity constraints (Liu et al. 2025b), and group causal transformers (Wang, Du, and Yu 2025) to model temporal visual changes. Yet, they overlook clinical context, limiting their capacity to generate personalized and context-aware reports. This gap leads to our central research question: Can we jointly model temporal visual changes and clinical context for improved cross-modal alignment and report generation?
@@
To address this challenge, we propose PriorRG, a novel chest X-ray report generation framework that mirrors real-world radiology workflows via a two-stage training pipeline. Stage 1 introduces a prior-guided contrastive pre-training scheme that simulates diagnostic reasoning by leveraging clinical context to guide spatiotemporal feature extraction, achieving better alignment with context-aware, disease progression-oriented radiology reports. Stage 2 presents a prior-aware coarse-to-fine decoding for report generation. We first devise an attention-enhanced layer fusion network to derive hierarchical visual representations from the vision encoder’s hidden states. Motivated by principles of visual cognition, PriorRG progressively integrates clinical context, spatiotemporal information, and hierarchical visual cues in a coarse-to-fine manner. This design equips the report generator with rich, context-aware, disease progression-oriented representations, thereby enhancing both the clinical efficacy and linguistic quality of the generated reports. Comprehensive experiments on the MIMIC-CXR (Johnson et al. 2019) and MIMIC-ABN (Ni et al. 2020) datasets demonstrate that our PriorRG significantly outperforms recent state-of-the-art methods in both medical image-text retrieval and radiology report generation. Our main contributions are:
@@ 2 Related Work
Radiology report generation (RRG). Unlike generic image captioning (Zeng et al. 2023; Hu and Li 2024), RRG aims to generate detailed clinical descriptions for medical images (Liu et al. 2025a). Recent advances have explored various techniques to improve accuracy, including knowledge graphs integration (Yin et al. 2025), contrastive learning (Liu et al. 2024b; Li et al. 2025b), retrieval-augmented methods (Jeong et al. 2024; Liu et al. 2024c), memory alignment (Chen et al. 2021; Shen et al. 2024), human preference optimization (Zhou et al. 2024b; Xiao et al. 2025), and LLM-based methods (Wang et al. 2023b; Liu et al. 2024a). However, most existing methods rely solely on single-view images and overlook patient-specific prior knowledge—an essential component in real-world clinical decision-making. This limitation hinders their ability to capture clinical intent and monitor disease progression. To address this gap, we propose PriorRG that integrates prior knowledge via coarse-to-fine decoding to generate context-aware, disease progression-oriented reports.
@@
RRG via prior knowledge. Several studies (Nguyen et al. 2023; Liu et al. 2024b; Mei et al. 2025) leverage indications to generate personalized reports, yet overlook disease progression—often leading to hallucinated descriptions of lesion changes. Recent efforts introduce longitudinal modeling via intra-modality similarity constraints (Liu et al. 2025b), report pre-filling (Zhu et al. 2023), or the group causal transformer (Wang, Du, and Yu 2025) to capture temporal changes. While promising, they still underexploit individual clinical context, limiting their ability to infer patients’ diagnostic intent. To bridge these gaps, we propose a prior-aware coarse-to-fine decoding that progressively integrates clinical context, disease progression patterns, and hierarchical visual cues, resulting in more accurate, fluent, and context-aware radiology reports.
@@
Figure 2: Overview of our PriorRG with a two-stage training pipeline, which consists of a vision encoder (RAD-DINO (Pérez-García et al. 2024)), a text encoder (CXR-BERT (Boecking et al. 2022)), and a report generator (DistilGPT2 (Sanh et al. 2020)).
@@
Medical vision-language pre-training (MVLP). MVLP aims to learn joint visual-textual representations to support downstream tasks such as image-text retrieval (Wang et al. 2022b; Zhang et al. 2023a; Bannur et al. 2023) and report generation (Jin et al. 2024; Liu et al. 2024b, 2025a). GLoRIA (Huang et al. 2021) and MGCA (Wang et al. 2022a) enhance representations using multi-level cross-modal alignment, while SEI (Liu et al. 2024b) links images to structured clinical entities. MedCLIP (Wang et al. 2022b) scales training via decoupled image-text matching based on the semantic similarity. To integrate domain knowledge, ARL (Chen, Li, and Wan 2022) and KAD (Zhang et al. 2023b) utilize UMLS (Bodenreider 2004) for semantic alignment and zero-shot disease classification. Beyond static image-text pairs, BioViL-T (Bannur et al. 2023) and MLRG (Liu et al. 2025a) leverage longitudinal data to model disease progression. However, clinical context—crucial for diagnostic reasoning—remains largely underexplored. To address this, we introduce a prior-guided contrastive pre-training scheme that explicitly encodes clinical context and tracks disease progression, thereby improving cross-modal alignment.
@@ 3.2 Stage 1: Prior-Guided Contrastive Pre-training
Visual features extraction. Following (Liu et al. 2025a), we employ the pre-trained vision encoder RAD-DINO (Pérez-García et al. 2024) to extract visual features from chest X-rays. Radiographic view positions (e.g., AP vs. PA) significantly affect image appearance—for instance, the cardiac silhouette appears enlarged in AP views. To account for such projection-specific variations, we introduce a learnable view-position embedding and fuse it with the extracted visual features, enhancing the model’s robustness to view-dependent discrepancies. The augmented features are then projected into a unified $d$-dimensional embedding space via a projection head. We denote the resulting visual features as ${\boldsymbol{V}}\in{\mathbb{R}^{M\times s\times d}}$, where $s$ is the sequence length, and $B$ represents batch size.
Textual features extraction. We adopt the pre-trained text encoder CXR-BERT (Boecking et al. 2022), followed by a projection head. Specifically, the last hidden state from CXR-BERT is passed through the projection head to obtain textual features, denoted as ${\boldsymbol{T}}\in{\mathbb{R}^{B\times p\times d}}$, where $p$ is the number of tokens. To ensure consistency across different types of input, we prepend special tokens—“[INDICATION]”, “[HISTORY]”, and “[FINDINGS]”—to the indication, medical history, and radiology reports, respectively (see Figure 2). This design facilitates type-aware feature extraction while gracefully handling missing fields (e.g., absent indication or medical history), enabling a unified and robust encoding process.
Spatiotemporal fusion network (STF). While the current image alone can support report generation, relying solely on it may lead to hallucinations—particularly producing disease progression description (e.g., “As compared to the previous radiograph, the patient has received a new right internal jugular vein catheter.”). To mitigate this, we employ a ViT-style spatiotemporal fusion network (Liu et al. 2025a; Dosovitskiy et al. 2021) to model disease progression between current and prior images. The fusion process proceeds as follows: First, we inject temporal embeddings into both current and prior visual features to encode chronological relationships. Each STF block is then formulated as:
@@
Instance-wise cross-modal alignment. In clinical practice, physicians typically begin with an initial clinical assessment based on a patient’s symptoms (i.e., indications) and medical history. Radiologists then integrate this clinical context with prior and current images to comprehensively evaluate lesion characteristics and progression. To simulate this diagnostic workflow, we incrementally incorporate clinical context ${\boldsymbol{T}}^{c}$ and spatiotemporal visual features ${\boldsymbol{\bar{V}}}^{st}$ using the Perceiver architecture (Jaegle et al. 2021), formulated as:
@@
To ensure the consistency between image-report pairs, we employ the instance-wise cross-modal alignment (Wang et al. 2022a; Liu et al. 2025a) to enhance multimodal representations. More concretely, we concatenate ${\boldsymbol{\bar{T}}}^{c}$ and ${\boldsymbol{\bar{V}}}^{st}$ along the sequence dimension, apply global average pooling, and perform L2 normalization to obtain global visual features ${\boldsymbol{V}}^{g}\in{\mathbb{R}^{B\times d}}$. The image-to-report similarity logits ${\boldsymbol{p}}^{i2r}\in{\mathbb{R}^{B\times B}}$ are then computed as:
@@ B-1 B-2 B-3 B-4 MTR R-L P R F1
Figure 3: Overview of attention-enhanced layer fusion network (ALF), which leverages CBAM (Woo et al. 2018) to extract hierarchical visual features.
@@ 3.3 Stage 2: Prior-Aware Coarse-to-Fine Decoding for Report Generation
Attention-enhanced layer fusion network. Existing methods (Liang et al. 2024; Liu et al. 2025a) utilize only the vision encoder’s last hidden state for report generation, which may neglect low-level details such as lesion morphology. To address this issue, we devise an attention-enhanced feature fusion network based on CBAM (Woo et al. 2018) (see Figure 3). Specifically, we apply CBAM’s channel attention and spatial attention modules to the hidden state of each encoder layer, dynamically reweighting features to emphasize diagnostically relevant information. The reweighted features are then processed by a channel fusion projector—implemented with Conv2D layers—which compresses and transforms the representations while retaining pixel-level and semantic information. Finally, the fused features are passed through a projection head, followed by temporal positional embedding and layer normalization, yielding hierarchical visual representations ${\boldsymbol{{V}}}^{hier}\in{\mathbb{R}^{B\times s\times d}}$ that capture both low-level details and high-level semantic cues.
Prior-aware coarse-to-fine decoding. Motivated by principles of visual cognition and radiologists’ diagnostic workflow, we propose a prior-aware coarse-to-fine decoding. Specifically, we first extract the condensed clinical context feature ${\boldsymbol{\bar{T}}}^{c}$ and the clinically informed spatiotemporal feature ${\boldsymbol{\bar{V}}}^{st}$ via Equations (3) and (4). These features are derived from the last hidden state of the text and vision encoders, and thus encode high-level semantics of patient clinical background and disease progression. We treat them as coarse-grained priors that offer a holistic overview for guiding subsequent fine-level decoding. We then enhance ${\boldsymbol{\bar{V}}}^{st}$ with hierarchical visual representations ${\boldsymbol{{V}}}^{hier}$ using the Perceiver architecture (Jaegle et al. 2021), formulated as:
@@
Report generation. We adopt the pre-trained DistilGPT2 (Sanh et al. 2020) to generate free-text reports. Stage 2 is trained to minimize the cross-entropy loss between the AI-generated report $\hat{y}_{i}$} and the radiologist-generated report $y_{i}$}. The objective function is formulated as:

### v2

@@ Abstract
Chest X-ray report generation aims to reduce radiologists’ workload by automatically producing high-quality preliminary reports. A critical yet underexplored aspect of this task is the effective use of patient-specific prior knowledge—including clinical context (e.g., symptoms, medical history) and the most recent prior image—which radiologists routinely rely on for diagnostic reasoning. Most existing methods generate reports from single images, neglecting this essential prior information and thus failing to capture diagnostic intent or disease progression. To bridge this gap, we propose PriorRG, a novel chest X-ray report generation framework that emulates real-world clinical workflows via a two-stage training pipeline. In Stage 1, we introduce a prior-guided contrastive pre-training scheme that leverages clinical context to guide spatiotemporal feature extraction, allowing the model to align more closely with the intrinsic spatiotemporal semantics in radiology reports. In Stage 2, we present a prior-aware coarse-to-fine decoding for report generation that progressively integrates patient-specific prior knowledge with the vision encoder’s hidden states. This decoding allows the model to align with diagnostic focus and track disease progression, thereby enhancing the clinical accuracy and fluency of the generated reports. Extensive experiments on MIMIC-CXR and MIMIC-ABN datasets demonstrate that PriorRG outperforms state-of-the-art methods, achieving a 3.6% BLEU-4 and 3.8% F1 score improvement on MIMIC-CXR, and a 5.9% BLEU-1 gain on MIMIC-ABN.
Codehttps://github.com/mk-runner/PriorRG
Extended versionhttps://arxiv.org/abs/2508.05353
@@ 1 Introduction
Figure 1: (a-c) show the workflow of existing methods and our approach. (d) summarizes their key properties. Longitudinal data includes both current and prior images. Patient-specific prior knowledge comprises the prior image (PI), indication, and medical history, which may be partially missing. “CC” and “DP” indicate whether a method models clinical context and disease progression, respectively.
@@
Radiology report generation (RRG) (aaai-2025-llm-rg4; 2024-acmmm-sptio-temporal-fusion) leverages AI techniques to automatically interpret medical images—such as chest X-rays (evoke; libra), CT scans (hamamci2024ct2rep; 2025-ct-rrg-nature-communications; MicarVLMoE-2025), and pathological slides (guo2024histgen; 2024-wsicaption-miccai)—and generate structured textual descriptions of clinically relevant findings. This automation supports radiologists by providing standardized preliminary reports, thereby improving diagnostic efficiency and consistency.
@@
While recent RRG approaches have achieved remarkable progress, most models operate on isolated images (see Figure 1(a)) and overlook essential patient-specific prior knowledge, including both clinical context (i.e., the indication and history sections) and the most recent prior image. Such prior knowledge plays a critical role in real-world clinical reasoning by enabling personalized interpretation and tracking of disease progression. However, existing methods (li-dcl; aaai-liu2024bootstrapping-llm; aaai-2025-mpo) largely ignore these factors, limiting their ability to produce context-aware, disease progression-oriented reports.
@@
To capture the clinical context, (ml4h-indication-rg) treats indications as auxiliary input while ignoring the noise in the data. SEI (sei) eliminates invalid or corrupted characters via preprocessing and employs a cross-modal fusion network to incorporate indications, thereby generating accurate findings. Similarly, for (2025-tmm-atl-ca). However, they fail to consider longitudinal information, often resulting in hallucinations when describing disease progression. To combat this issue, existing methods incorporate prior images (Figure 1(b)) and employ techniques such as report pre-filling (pre-fill-longitudinal-miccai-2023), intra-modality similarity constraints (aaai-2025-hc-llm), and group causal transformers (2024-eccv-hergen) to model temporal visual changes. Yet, they overlook clinical context, limiting their capacity to generate personalized and context-aware reports. This gap leads to our central research question: Can we jointly model temporal visual changes and clinical context for improved cross-modal alignment and report generation?
@@
To address this challenge, we propose PriorRG, a novel chest X-ray report generation framework that mirrors real-world radiology workflows via a two-stage training pipeline. Stage 1 introduces a prior-guided contrastive pre-training scheme that simulates diagnostic reasoning by leveraging clinical context to guide spatiotemporal feature extraction, achieving better alignment with context-aware, disease progression-oriented radiology reports. Stage 2 presents a prior-aware coarse-to-fine decoding for report generation. We first devise an attention-enhanced layer fusion network to derive hierarchical visual representations from the vision encoder’s hidden states. Motivated by principles of visual cognition, PriorRG progressively integrates clinical context, spatiotemporal information, and hierarchical visual cues in a coarse-to-fine manner. This design equips the report generator with rich, context-aware, disease progression-oriented representations, thereby enhancing both the clinical efficacy and linguistic quality of the generated reports. Comprehensive experiments on the MIMIC-CXR (johnson-mimic-cxr-jpg) and MIMIC-ABN (mimic-abn-ori) datasets demonstrate that our PriorRG significantly outperforms recent state-of-the-art methods in both medical image-text retrieval and radiology report generation. Our main contributions are:
@@ 2 Related Work
Radiology report generation (RRG). Unlike generic image captioning (cvpr-2023conzic-image-captioning; 2024-acmmm-image-caption), RRG aims to generate detailed clinical descriptions for medical images (mlrg). Recent advances have explored various techniques to improve accuracy, including knowledge graphs integration (2025-kia-coling-knowledge), contrastive learning (sei; cofe-eccv-24), retrieval-augmented methods (jeong2024multimodal-MIDL; fse), memory alignment (chen-etal-2021-cross-modal; shen2024automatic_aaai), human preference optimization (2024-llm-cxr-rl; aaai-2025-mpo), and LLM-based methods (wang-2023-r2gengpt; aaai-liu2024bootstrapping-llm). However, most existing methods rely solely on single-view images and overlook patient-specific prior knowledge—an essential component in real-world clinical decision-making. This limitation hinders their ability to capture clinical intent and monitor disease progression. To address this gap, we propose PriorRG that integrates prior knowledge via coarse-to-fine decoding to generate context-aware, disease progression-oriented reports.
@@
RRG via prior knowledge. Several studies (ml4h-indication-rg; sei; 2025-tmm-atl-ca) leverage indications to generate personalized reports, yet overlook disease progression—often leading to hallucinated descriptions of lesion changes. Recent efforts introduce longitudinal modeling via intra-modality similarity constraints (aaai-2025-hc-llm), report pre-filling (pre-fill-longitudinal-miccai-2023), or the group causal transformer (2024-eccv-hergen) to capture temporal changes. While promising, they still underexploit individual clinical context, limiting their ability to infer patients’ diagnostic intent. To bridge these gaps, we propose a prior-aware coarse-to-fine decoding that progressively integrates clinical context, disease progression patterns, and hierarchical visual cues, resulting in more accurate, fluent, and context-aware radiology reports.
@@
Figure 2: Overview of our PriorRG with a two-stage training pipeline, which consists of a vision encoder (RAD-DINO (2024-rad-dino)), a text encoder (CXR-BERT (2022-eccv-cxr-bert)), and a report generator (DistilGPT2 (Sanh2019DistilBERTAD)).
@@
Medical vision-language pre-training (MVLP). MVLP aims to learn joint visual-textual representations to support downstream tasks such as image-text retrieval (medclip_wang_2022; zhang-2023-biomedclip; 2023-cvpr-biovit-2301) and report generation (promtmrg-aaai-2024; sei; mlrg). GLoRIA (Huang_2021_ICCV_gloria) and MGCA (wang-mgca) enhance representations using multi-level cross-modal alignment, while SEI (sei) links images to structured clinical entities. MedCLIP (medclip_wang_2022) scales training via decoupled image-text matching based on the semantic similarity. To integrate domain knowledge, ARL (chen2022align) and KAD (zhang-kad) utilize UMLS for semantic alignment and zero-shot disease classification. Beyond static image-text pairs, BioViL-T (2023-cvpr-biovit-2301) and MLRG (mlrg) leverage longitudinal data to model disease progression. However, clinical context—crucial for diagnostic reasoning—remains largely underexplored. To address this, we introduce a prior-guided contrastive pre-training scheme that explicitly encodes clinical context and tracks disease progression, thereby improving cross-modal alignment.
@@ 3.2 Stage 1: Prior-Guided Contrastive Pre-training
Visual features extraction. Following (mlrg), we employ the pre-trained vision encoder RAD-DINO (2024-rad-dino) to extract visual features from chest X-rays. Radiographic view positions (e.g., AP vs. PA) significantly affect image appearance—for instance, the cardiac silhouette appears enlarged in AP views. To account for such projection-specific variations, we introduce a learnable view-position embedding and fuse it with the extracted visual features, enhancing the model’s robustness to view-dependent discrepancies. The augmented features are then projected into a unified $d$-dimensional embedding space via a projection head. We denote the resulting visual features as ${\boldsymbol{V}}\in{\mathbb{R}^{M\times s\times d}}$, where $s$ is the sequence length, and $B$ represents batch size.
Textual features extraction. We adopt the pre-trained text encoder CXR-BERT (2022-eccv-cxr-bert), followed by a projection head. Specifically, the last hidden state from CXR-BERT is passed through the projection head to obtain textual features, denoted as ${\boldsymbol{T}}\in{\mathbb{R}^{B\times p\times d}}$, where $p$ is the number of tokens. To ensure consistency across different types of input, we prepend special tokens—“[INDICATION]”, “[HISTORY]”, and “[FINDINGS]”—to the indication, medical history, and radiology reports, respectively (see Figure 2). This design facilitates type-aware feature extraction while gracefully handling missing fields (e.g., absent indication or medical history), enabling a unified and robust encoding process.
Spatiotemporal fusion network (STF). While the current image alone can support report generation, relying solely on it may lead to hallucinations—particularly producing disease progression description (e.g., “As compared to the previous radiograph, the patient has received a new right internal jugular vein catheter.”). To mitigate this, we employ a ViT-style spatiotemporal fusion network (mlrg; 2021-vit) to model disease progression between current and prior images. The fusion process proceeds as follows: First, we inject temporal embeddings into both current and prior visual features to encode chronological relationships. Each STF block is then formulated as:
@@
Instance-wise cross-modal alignment. In clinical practice, physicians typically begin with an initial clinical assessment based on a patient’s symptoms (i.e., indications) and medical history. Radiologists then integrate this clinical context with prior and current images to comprehensively evaluate lesion characteristics and progression. To simulate this diagnostic workflow, we incrementally incorporate clinical context ${\boldsymbol{T}}^{c}$ and spatiotemporal visual features ${\boldsymbol{\bar{V}}}^{st}$ using the Perceiver architecture (icml-2021-perceiver), formulated as:
@@
To ensure the consistency between image-report pairs, we employ the instance-wise cross-modal alignment (wang-mgca; mlrg) to enhance multimodal representations. More concretely, we concatenate ${\boldsymbol{\bar{T}}}^{c}$ and ${\boldsymbol{\bar{V}}}^{st}$ along the sequence dimension, apply global average pooling, and perform L2 normalization to obtain global visual features ${\boldsymbol{V}}^{g}\in{\mathbb{R}^{B\times d}}$. The image-to-report similarity logits ${\boldsymbol{p}}^{i2r}\in{\mathbb{R}^{B\times B}}$ are then computed as:
@@ B-1 B-2 B-3 B-4 MTR R-L P R F1
Figure 3: Overview of attention-enhanced layer fusion network (ALF), which leverages CBAM (Woo_2018_ECCV-cbam) to extract hierarchical visual features.
@@ 3.3 Stage 2: Prior-Aware Coarse-to-Fine Decoding for Report Generation
Attention-enhanced layer fusion network. Previous methods (liang2024divide-acmmm-24; mlrg) rely solely on the vision encoder’s last hidden state, overlooking low-level details such as lesion morphology. To mitigate this, we design an attention-enhanced layer fusion network based on CBAM (Woo_2018_ECCV-cbam) (see Figure 3). CBAM’s channel and spatial attention are applied to each encoder layer to highlight diagnostically relevant features. The refined features are fused through a Conv2D projector that preserves both spatial and semantic information. A projection head with temporal positional embedding and layer normalization then yields hierarchical visual representations ${\boldsymbol{V}}^{hier}\in\mathbb{R}^{B\times s\times d}$ capturing multi-level cues.
Prior-aware coarse-to-fine decoding. Motivated by principles of visual cognition and radiologists’ diagnostic workflow, we propose a prior-aware coarse-to-fine decoding. Specifically, we first extract the condensed clinical context feature ${\boldsymbol{\bar{T}}}^{c}$ and the clinically informed spatiotemporal feature ${\boldsymbol{\bar{V}}}^{st}$ via Equations (3) and (4). These features are derived from the last hidden state of the text and vision encoders, and thus encode high-level semantics of patient clinical background and disease progression. We treat them as coarse-grained priors that offer a holistic overview for guiding subsequent fine-level decoding. We then enhance ${\boldsymbol{\bar{V}}}^{st}$ with hierarchical visual representations ${\boldsymbol{{V}}}^{hier}$ using the Perceiver architecture (icml-2021-perceiver), formulated as:
@@
Report generation. We adopt the pre-trained DistilGPT2 (Sanh2019DistilBERTAD) to generate free-text reports. Stage 2 is trained to minimize the cross-entropy loss between the AI-generated report $\hat{y}_{i}$} and the radiologist-generated report $y_{i}$}. The objective function is formulated as: